The information in the folowing two way contingency table relates the make of a new light vehicle (American, Asian, European) and the age of the person
18-34 35-64 65 and older
American 0.06 0.31 0.10
Asian 0.05 0.30 0.07
European 0.01 0.05 0.05
A) The probability that a randomly seleted new light vehicle is a European make is about?
B) The probability that a randomly seleted new light vehicle is a European make, given that the purchaser is 65 and older, is about?

Answer by robertb(5830)

A) p(European) = 0.01+0.05+0.05 = 0.11
B) p(European|65 or older) = p(European & 65 and older)/p(65 and older) = 0.05/(0.10+0.07+0.05) = 0.05/0.22 = 5/22.